OpenStudy (anonymous):

Find an angle between 0 and 2pie(3.14) that is coterminal with -13/24pie(3.14)

OpenStudy (anonymous):

\[48/24* \pi - 13/24 * \pi = 35/24* \pi \]

OpenStudy (anonymous):

Although the normal formula is: \[ \theta - 2 \pi \] and \[ \theta + 2 \pi\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Np.